OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by the equation below. y = x^2 + x - 12 A. (-12, 0) and (1, 0) B. (-4, 0) and (3, 0) C. (0, -4) and (0, 3) D. (0, -12) and (0, 1)

OpenStudy (anonymous):

get the equation into factored form like: you need to think about what two numbers can you multiply together and get -12 and then add together to get +1

OpenStudy (anonymous):

(x-3)(x+4) set each expression equal to 0 and solve. (x-3) =0 (x+4) =0

OpenStudy (anonymous):

remember that an x intercept is where the line crosses the x axis which means it will be wherever the y value is 0

OpenStudy (anonymous):

@ehuman why not let him think about the problem instead of blurting out the answer?

OpenStudy (anonymous):

the answer is not blurted nor is it written, the first step was given they still have to work towards the solution